La présentation est en train de télécharger. S'il vous plaît, attendez

La présentation est en train de télécharger. S'il vous plaît, attendez

SPS LS1 Planning meeting (N011) 10/04/13

Présentations similaires


Présentation au sujet: "SPS LS1 Planning meeting (N011) 10/04/13"— Transcription de la présentation:

1 SPS LS1 Planning meeting (N011) 10/04/13
David Mcfarlane

2 Topics Key dates at the start of LS1. Lifts and Monte-charges.
Test AUGs and bank holidays. Consignations. Impacts. Removal of magnets from BA5. LS1 progress report. Planning. False floors lifted in Surface Bas. E.C.Rs. Important Information. A.O.B

3 Key dates at the start of LS1
Description Date HV Tests Monday the 25th March 2013 GS/ASE will switch each access point over to “General Access” mode Tuesday the 26th March 2013 General access into the SPS/BA7/TI2/TI8/TT10 allowed Wednesday the 27th March 2013 Technical network disconnection test 27th March 2013 (10:00 to 13:00) Controls maintenance 2nd April to 4th April 2013 (TBC) LIMITED Access into BA80 25th September 2013

4 Lift Maintenance March / April
Pour raisons de sécurité , le monte-charge 25 Tonnes du BA7 AS513 est mis à l’arrêt en attendant de faire le test de charge. (Test prévu semaine 16)

5 Test AUG/ Bank holidays (Access will be blocked)
Location Date (AUG tests) BA4 9th April (AUG tests) BA5 Wednesday the 10th April (AUG tests) BA7 Thursday the 11th April May day Wednesday the 1st May (AUG) BA6 Tuesday the 7th May Ascension Thursday the 9th May Whit Monday Monday the 20th May Les maintenance des tableaux électrique EBD% des BA1 – BA3 – BA4 – BA6 – BA7 – se dérouleront le même jour que les TEST AUG. Ceci pour ne pas couper l’électricité une seconde fois cette année sur chaque point du SPS. Les coupures dureront donc de 7h00 à 22h00 lors des test AUG et maintenance. Les notes de coupures suivront et seront diffusées.

6 Consignations Please ensure that The coordination team has a copy of all main consignations of the SPS. It is the responsibility of the Equipment owner to consign any equipment that they will be working on. Do not assume something is consigned. Check first.

7 Discipline / Spécialité
“Standard” IMPACTs Discipline / Spécialité Description Commentaires No. Impact Responsable BE/ABP IMPACT STANDARD: campagnes générales de relevés géométriques 24163 Patrick Bestmann BE/BI IMPACT STANDARD: maintenances générales de l’instrumentation 24203 Stephan Burger BE/CO IMPACT STANDARD: maintenances générales syst. contrôle 24204 Claude Dehavay BE/RF IMPACT STANDARD: maintenances générales syst. RF 24206 Christophe Renaud EN/CV IMPACT STANDARD: maintenance générale syst refr. et ventilation 24207 Bill Bannister EN/EL IMPACT STANDARD: interventions électriques générales (câbles, connecteurs, éclairage, coffrets, …) 24208 Guillaume Gros EN/EL/CF IMPACT STANDARD: interventions fibre optics 24209 Simao Pedro Costa Machado EN/EL/OP IMPACT STANDARD: interventions Opération Réseaux 24210 Denis Ribiollet EN/HE - transport IMPACT STANDARD: transports – manutentions 24212 Patrick Vallet EN/HE - maintenance IMPACT STANDARD: maintenances générales des équipements de transport, levage et ascenseurs 24202 Nicolas Perez

8 Discipline / Spécialité
“Standard” IMPACTs Discipline / Spécialité Description Commentaires No. Impact Responsable EN/STI - stoppers IMPACT STANDARD: maintenances générales beam stoppers 24213 Jerome Lendaro EN/STI – UA9 IMPACT STANDARD: maintenances générales UA9 Experience 24214 GS/ASE IMPACT STANDARD: interventions générales syst. accès, alarme, détection incendie 24216 Didier Vaxelaire / Denis Raffourt GS/SE IMPACT STANDARD: maintenances et réparation ouvrages souterrains 24217 Richard Morton TE/ABT - kickers IMPACT STANDARD: maintenances générales kickers 24218 Francesco Castronuovo TE/ABT - septa IMPACT STANDARD: maintenances générales septa 24219 Cedric Baud TE/MSC IMPACT STANDARD: maintenance générale sur les aimants « magnet patrols » et tests HV feront plutôt l’objet d’un Impact spécifique 24222 Jeremie Bauche TE/VSC IMPACT STANDARD: maintenance générale sur les équipements Vide 24224 Jose A. Ferreira TE/EPC IMPACT STANDARD: maintenance transformers 24226 Gilles Le Godec Visites et Inspections - Coordination IMPACT STANDARD: coordination générale, suivi des activités, vérification des plans de layout, … essentiellement groupe EN/MEF 24227 David Mcfarlane / Franck Bais / Frederic Galleazzi

9 Discipline / Spécialité
“Standard” IMPACTs Discipline / Spécialité Description Commentaires No. Impact Responsable Visites et Inspections - RP IMPACT STANDARD: contrôles RP, suivi des activités par RP essentiellement section DGS/RP/AS 24229 Christophe Tromel / Angelito Herve Visites et Inspections - OP IMPACT STANDARD: patrouilles, tests DSO, réarmements portes, … essentiellement groupe BE/OP 24230 James Ridewood Visites et Inspections - Sécurité IMPACT STANDARD: inspections sécurité, VIC, pompiers, gardiennage… 24231 Emmanuel Paulat / John Etheridge Visites et Inspections - Visiteurs IMPACT STANDARD: visiteurs externes, visites professionnelles, … 24233 Franck Bais / Marc Tavlet / Steve Hutchins / John Pedersen RP piquet Piquet radioprotection 24334 C. Tromel EN/EL/BT IMPACT STANDARD: Maintenance électriques générales (transformateurs, 48V, éclairage, coffrets, tableaux basse tension) 24506 Stephane Delarue TE/MPE IMPACT STANDARD: maintenances générales des équipements de TE/MPE 24885 Richard Mompo IS/CS IT/CS piquet 25005 Gyorgy Balazs TE/HDA Safety Officer Patrols in the SPS 24508 Thomas Otto

10 “Specific Task” IMPACTs
IMPACTs are being created for specific tasks. (eg. Changing of a Kicker magnet) This IMPACT will be created by the primary team responsible for the work. All teams involved with this task must ensure that they are included on this IMPACT. Tasks with Specific IMPACTs are indicated on the planning schedule. (BLUE)

11 Removal of coated magnets from BA5
Zone Position Secteur vide Type aimant N° aimant Ordre de remplacement Date Heure de sortie (approximative) Lieu de destockage 5- 51310 510 QD Q 136 1 vendredi 5 avril 10h00 selon contrôle RP 867 ou 954 51290 MBB MB 261 2 12h00 51270 MB 142 3 14h00 51250 MBA MA 245 4 16h00 51230 MA 124 5 Lundi 8 avril 51190 MA 204 6 51170 MA 169 7 51150 MB 258 8 51130 MB 242 9 Jeudi 11 avril 51090 MB 255 10 51070 MB 252 11 51050 MA 277 12 51030 MA 75 13 Vendredi 12 avril 51010 QF Q 125 14 51210 Q 174 15 51110 Q 122 16

12 LS1 progress report An Excel Progress report has been created to present to management the status of progress compared to the planning for LS1.

13 SPS LS1 schedule Beam on EPC Commissioning tests
Beam off : 16/02/13 Beam on: 13/10/14 Open Access : 27/03/13 Close Access : 27/06/14 Beam on Cool down/RP survey and Magnet tests Cool down NO ACCESS Fibres 18 KV Kicker Survey Civil Engineering in TT10 RF upgrade works Survey RF upgrade works EPC works Cables Survey Survey Kicker Cables Survey EPC works Survey Survey Maintenance General Maintenance General Magnets Magnets Magnets Magnets Kicker Magnets Magnets Cables Magnets 18 KV Magnets General maintenance General maintenance Crab Cavities General Maintenance Cooling tower 863 General Maintenance Survey Vacuum Ring-main Ring-main Ring-main Vacuum Ring-main Septa works Ring-main Vacuum Ring-main Fibres Septa works Ring-main Prep work in LSS1 18 KV General Maintenance Maintenance General Vacuum Vacuum RF Upgrade works Cables Survey Horizontal Vertical smoothing Cool down Limited ACCESS Fibres TT20 pipework 18 KV CV maintenance CV maintenance UA works Civil Engineering Irradiated Cables 18 KV Fibres CV maintenance Fibres Cables Cables Cables 18 KV Cables EL Studies Civil Engineering TT10 Consolidation Cables replacement Septa Civil Engineering Civil Engineering CV studies RF Upgrade works Crab Cavities Civil Engineering Civil Engineering RF Upgrade works Re-install equipment removed for cable works RAL Magnets CV works Fibres Cables Magnets Magnets Magnets Magnets Magnets Magnets Magnets EPC Commissioning tests Vacuum Cold check out & Start-up DSO tests/HV tests Protons to the North Area Protons to LHC Ions to NA61 CV Controls upgrade Re-route Cabling

14 False floors lifted in Surface BAs
Below is a table showing the dates that EL will be pulling cables at each point of the SPS. During these periods large sections of the false floor in these surface buildings will be lifted. Any works in the BA’s that involve lifting of the false floors MUST be reported to the Coordination team BAs Cabling pulling start Cable pulling end BA1 (TS1- works) BA1 (Irradiated cable works) BA1 (Fibre Optics) 29/04/2013 08/10/2013 08/04/2013 02/08/2013 14/02/2014 24/04/2013 BA2 06/01/2014 11/03/2014 BA3 BA3 (Fibre Optics) 27/03/2013 22/04/2014 31/05/2013 06/05/2014 BA4 01/07/2013 16/08/2013 BA5 BA5 (Fibre Optics) 30/09/2013 24/05/3013 11/10/2013 BA6 BA6 (Fibre Optics) 13/15/2013 18/10/2013 24/05/2013 BA7 16/10/2013 07/11/2013 If you require access to racks or other equipment in the surface buildings during these periods you must check with planning first.

15 Engineering Change Requests
ECR link ECR Procedure ECR Template All ECRs must be submitted to Jerome Axensalva (BE/OP) AND myself.

16 Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 Removal of AEWB Pick-up from position to 31502 (EDMS No ) RELEASED Removal of a BPWA Pick-up from position to 31101 (EDMS No ) RELEASED Installation of a new Wideband Kickers system in position 32173 (EDMS No ) RELEASED Installation of a new HBW Damper Pick-up in place of BPWA & AEWB 31932 (EDMS No ) RELEASED Replacement of 4 old BPA and 2 BPD by 6 BPCE in LSS1 & LSS2 (EDMS No ) RELEASED Reinstallation of pump and gauges (Pirani + Penning) in (vacuum sector 150) (EDMS No ) RELEASED Modification LSS1 sectorization (new VVSF + displacement VVSB 11772) (EDMS No v.2) UNDER APPROVAL Installation of 3 BCT in LSS5 for North Area protection during Ions operations in SPS. (EDMS No ) UNDER IMPLEMENTATION

17 Current ECR’s for LS1 https://edms.cern.ch/nav/P:CERN :V0/P:CERN :V0/TAB3 Transfer of the Fast Beam Current Transformers (BCTFR.31451) for the SPS circulating beam from LSS3 to LSS5. (EDMS No ) RELEASED SPS Beam Position Monitoring Prototyping Setup in LSS4. (EDMS No ) RELEASED Installation de 4 moniteurs de position (2 BPH, 2 BPV), dédiés au système Damper. (EDMS No ) RELEASED Installation d'un nouveau moniteur de position de type BPCR en LSS2 au (EDMS No ) RELEASED Displacement of Door PPG2145 separating the LHC and the SPS in the TI2 tunnel. (EDMS No ) RELEASED Interlocking of Injection Tunnel upstream areas with LHC powering. (EDMS No ) RELEASED Move of BPH to BPH (EDMS No ) UNDER APPROVAL Optical Fibre Racks. (EDMS No ) RELEASED

18 Important information
Access requests Ensure everyone entering the SPS has the correct access authorisation (via EDH) and has completed the correct courses. The following safety equipment is Obligatory in the SPS for EVERYONE! Safety shoes Biocell Electronic dosimetre (Make sure you have your new one from Building 55!) Passive dosimetre Helmet plus Light Useful Contact Numbers SPS Coordination Team David MCFARLANE Franck BAIS – James Ridewood Control Room SPS CCC – 77500 TI CCC – 72201 CSA RP Christophe Tromel – Angelito Herve – Nadine Conan (BA1) – Kurt Weiss (BA1) Coordination Safety Emmanuel Paulat – John Etheridge

19 A.O.B. Tractor update Nous avons des tanks MSE a transporter du BA3 au BA6, Nous devons passer par le BA4 et le BA5 en date du mardi 16 Avril avec 2 tanks, déchargement au BA6 et retour 1 remorque vide au BA3. Le mercredi 17 Avril transport 1 tank du BA3 au BA6 , retour des tracteurs (2) au BA3. Nous restons en contact avec guillaume pour la meilleur organisation pour effectuer ces transfert. Moving of PPG2145 in TI2 See photos (Taken by Frederic Galleazzi) A.Herve signale que plusieurs personnes accèdent au BA1 n’importe quand (non respect du planning) et annonce la remise du VETO si cela continu.

20 SPS coordination website
\\cernhomeJ.cern.ch\J\jocksoft\Public\LS1 Next meeting 08/05/2013 892-1-D20 (Usual room)


Télécharger ppt "SPS LS1 Planning meeting (N011) 10/04/13"

Présentations similaires


Annonces Google